Hellium takes 5 seconds to effuse from a hole of 10dm3 containers. How long would it take for oxygen to effuse from the some container at the same temperature and pressure? 



1
Expert's answer
2022-01-11T23:43:01-0500

According to Graham's law of Effusion/diffusion


"r\propto\sqrt\frac{1}{M}"


Where "r=" Rate of effusion

"M=" Molar mass


Rate of effusion "=" "\frac{Volume}{time}"


"\frac{Rate\>of\>He}{Rate\>of\>O_2}=\frac{M_{O_2}}{M_{He}}"


"\frac{\frac{10}{5}}{\frac{10}{T}}=\sqrt\frac{32}{4}"


"T=14.14\>seconds"